OpenStudy (anonymous):

How would I set up this problem? - Minnie deposited $5,375 into a savings account 18 years ago. The account has an interest rate of 3.2% and the balance is currently $9,561.39. How often does the interest compound?

OpenStudy (tkhunny):

Solve! \(5375.00\left(1 + \dfrac{0.032}{m}\right)^{18*m} = 9561.39\) Of course, you cannot solve that for m, so I'd just try a few values. \(5375.00\left(1 + \dfrac{0.032}{1}\right)^{18*1} = ??\) \(5375.00\left(1 + \dfrac{0.032}{2}\right)^{18*2} = ??\) \(5375.00\left(1 + \dfrac{0.032}{4}\right)^{18*4} = ??\) \(5375.00\left(1 + \dfrac{0.032}{12}\right)^{18*12} = ??\) If none of those works, we'll just have to work a little harder. Note: Continuous compounding gives $9,561.50. Maybe we should try: \(5375.00\left(1 + \dfrac{0.032}{52}\right)^{18*52} = ??\) \(5375.00\left(1 + \dfrac{0.032}{365}\right)^{18*365} = ??\)